OpenStudy (anonymous):

A 15 foot ladder is resting against the wall. The bottom is initially 10 feet away from the wall and is being pushed towards the wall at a rate of 1/4 ft/sec. How fast is the top of the ladder moving up the wall 12 seconds after we start pushing?

OpenStudy (perl):



OpenStudy (perl):

x^2 + y^2 = z^2 2x * dx/dt + 2y*dy/dt = 2z * dz/dt divide out 2 z is constant, so dz/dt = 0 x*dx/dt + y * dy/dt = 0

OpenStudy (perl):

A 15 foot ladder is resting against the wall. The bottom is initially 10 feet away from the wall and is being pushed towards the wall at a rate of 1/4 ft/sec. How fast is the top of the ladder moving up the wall 12 seconds after we start pushing? so x = 10, dx/dt = -1/4 , and solve for dy/dt

OpenStudy (perl):

10*(-1/4) + y * dy/dt = 0 we can find y by using pythagorean theorem 10^2 + y^2 = 15^2